OpenStudy (nikita2):

Could you help me with this equation y'' = 1/y^2

OpenStudy (jamesj):

This is a tricky equation. I'll come back in a minute and help you with it.

OpenStudy (anonymous):

are u asking for y?

OpenStudy (shayaan_mustafa):

it is a differential equation.

OpenStudy (shayaan_mustafa):

ask me what are you trying to ask?

OpenStudy (nikita2):

yes it is differential eq.

OpenStudy (shayaan_mustafa):

so what are asking for? finding general solution or anything else?

OpenStudy (nikita2):

general

OpenStudy (shayaan_mustafa):

ok you need a general solution. There are two solution. first is complementary solution and other is particular solution. do you know this?

OpenStudy (shayaan_mustafa):

you just get relax nikita2. I am here for help. take time and post a correct and complete question.

OpenStudy (nikita2):

ok, I will )

OpenStudy (shayaan_mustafa):

ok

OpenStudy (jamesj):

@shayaan, do you know how to find the particular solution? This is quite non-standard.

OpenStudy (shayaan_mustafa):

yes i know. there are several methods. but good one is variation of parameters .

OpenStudy (jamesj):

No, that's not correct. This is non-linear equation.

OpenStudy (nikita2):

\[ (D _{2}^{t}u)^{2} (D _{2}^{x}D _{2}^{t}u) = 1\] And i should to solve Cauchy problem

OpenStudy (shayaan_mustafa):

is it non-linear?

OpenStudy (nikita2):

yes(

OpenStudy (shayaan_mustafa):

hmmm... then fine.

OpenStudy (jamesj):

y = y(t), y'' - 1/y^2 = 0. Not linear, unfortunately.

OpenStudy (shayaan_mustafa):

do you know how to start with cauchy euler method?? you get start then we are here to volunteerily

OpenStudy (nikita2):

yes i'll try it now

OpenStudy (jamesj):

ok ... multiply first both sides by y' dt \[ y'' y' dt = \frac{y'}{y^2} dt \] Writing \( d(y') = y' dt \) and \( dy = y' dt \) we have \[ (y')' d(y') = \frac{1}{y^2} dy \] Now integrate both sides and we have ....

OpenStudy (shayaan_mustafa):

@JamesJ give him reason why we did so? So that he could get clear concepts. Thanks.

OpenStudy (nikita2):

Now integrate both sides and we have .... \[(y')^2/2 = -1/3*1/y^3\]

OpenStudy (nikita2):

+ Const

OpenStudy (jamesj):

\[ \frac{1}{2} (y')^2 = -\frac{1}{y} + C \]

OpenStudy (nikita2):

Oh! i've done mistake here!

OpenStudy (jamesj):

Now.... \[ \frac{dy}{dt} = \pm \sqrt{ \frac{ 2(Cy-1)}{y}} \] This is separable and it isn't too hard to find an equation for t, t = t(y).

OpenStudy (nikita2):

i see...

OpenStudy (jamesj):

Interesting, eh? I only figured this out recently myself, working it out in this problem: http://openstudy.com/study#/updates/4f0de661e4b084a815fcff56

OpenStudy (nikita2):

Oh! I've not notise that it is forse of atraktion!

OpenStudy (nikita2):

But anyway if i'll try to do next. Solve \[D _{2}^{t}u(t,x) = y(t,x)\] where y we found

OpenStudy (jamesj):

Tricky, but probably not impossible analytically. What's the original motivation for this problem?

OpenStudy (nikita2):

I solving the Cauchy problem \[(D _{2}^{t}u(t,x))^2*(D _{2}^{x}D _{2}^{t}u(t,x)) =\] the initial condition is on {t=0}

OpenStudy (nikita2):

=1

OpenStudy (jamesj):

right. btw, you might find this site helpful: math.stackexchange.com

OpenStudy (jamesj):

but in the meantime, it would be great if you wanted to help out here answering some elementary questions now and again!

OpenStudy (nikita2):

I did it and I will )

OpenStudy (nikita2):

Thank you very much!

OpenStudy (jamesj):

Happy to help. This was a fun problem.